Reports 1-1 of 1 Clear search Modify search
AdV-SGD (FDS commissioning)
capocasa - 20:57 Monday 19 July 2021 (52603) Print this report
FC relocked -- lower injected green power

After opening of the FCIM gate vales, filter cavity has been realigned and relocked. The lock was quite stable but the maximum transmission is only 1.7 volt.

I went to check on EQB1 and I found that the green power after Faraday isolator is only 7.5 mW. All the iris to cut AOM high orders were found open. Moreover the beam looked as it is a bit clipped already before the FI.

=========================

Some sinusoidal noise injection on FCEM TX have been performed from gps 1310735174 for the dithering implementation.

 

Images attached to this report
Search Help
×

Warning

×